OpenStudy (anonymous):

solve equation 2x+5=3

OpenStudy (anonymous):

4

OpenStudy (theviper):

2x=3-5=-2 x=-2/2=-1

OpenStudy (anonymous):

in order to solve for x, isolate the term with x by subtracting _____ on ____ sides. then divide both sides by ____ to obtain x

OpenStudy (anonymous):

retype problem again

OpenStudy (anonymous):

rovg, delete your responses

OpenStudy (anonymous):

sorry didn't see the whole problem and got in too late srry

OpenStudy (anonymous):

2t+5=3

OpenStudy (theviper):

2t+5=3 2t=3-5=-2 t=-1

OpenStudy (anonymous):

2t+5=3 -5 -5 2t=-2 /2 /2 t=-1

OpenStudy (anonymous):

why do it like this 2t+5=3 2t=3-5 =-2 so therefore 2t=-2 divde by 2/-2 t=-1